OpenStudy (anonymous):

Can someone explain some SAT math questions? --- 1. If x^2 + y^2 = 30 and xy = 8. What is the value of (x-y)^2 The answer is 14 --- 2. If 2^(3n+2) = 4^(n+4), what is the value of n. The answer is 6. --- 3. If k and h are constants and x^2 - kx + 18 = (x-3)(x-h), what is the value of k. The answer is 9. ---

OpenStudy (hari5719):

sat is hard

OpenStudy (anonymous):

I will only give you an example of SAT.

OpenStudy (anonymous):

What is the area, in square feet, of the triangle whose sides have lengths equal to 10, 6 and 8 feet? Solution If given the three sides of a triangle and asked to find the area of the triangle, we normally use Heron's formula. However the triangle given here is a special triangle. Note that the three sides of the given triangle are related as follows √(62 + 82) = 10 which means that the triangle is a right triangle and its hypotenuse is 10 and legs 6 and 8. The area A is given by A = (1/2)(6)(8) = 24 feet2

OpenStudy (anonymous):

Sorry, but that doesn't help me solve algebra problems. >.< I took the test which had those questions, but don't understand them.

OpenStudy (anonymous):

algebra? are you in a grade schooler student?

OpenStudy (anonymous):

No, it's just SAT prep.

OpenStudy (anonymous):

x+2y=1......(1) 2x+y=8......(2) We will use the elimination method. First, multiply (1) by -2 and add to (2) -2x-4y=-2 2x+y=8 ==> -3y=6 ==> y=-2 x+2y=1 ==> x= 1-2y = 1-2(-2)= 1+4= 5 like this ? what do you think?

hartnn (hartnn):

hints : 1) \(\Large (x-y)^2 = x^2 +y^2 -2xy\) 2) write 4 = 2^2, then compare the exponents 3) FOIL, right side, then compare the co-efficients